# Dehua Tao1, Daxin Tan2, Yu Ting Yeung2, Xiao Chen2, Tan Lee1

### v1

@@ 5.2 Ablation study
A con of ToneUnit is the requirement of annotated data with tone labels. To evaluate the required amount of labeled data, we perform ablation under the same configuration as the FSQ model (ID 4) with different sizes of AISHELL-1 training set. We randomly sample 3 training subsets of 1, 10, and 50 hours from the full 150-hour training set for the fine-tuning process. We still train the speech synthesizer with the full 150-hour AISHELL-1 dataset for 100k updates. Note that the speech synthesizer is trained with extracted discrete speech units without any text labels. The results are reported in Table 3 in terms of codebook usage and CER. FSQ can still fully use the codewords at different sizes of labeled data. With just 1 hour of annotated data with tone labels, the ToneUnit framework successfully generates speech units capable of synthesizing Mandarin Chinese speech at acceptable machine intelligibility. With a subset of 10-hour training data, CER of synthesized speech is already comparable to those from the full training set. This suggests that the proposed ToneUnit framework retains effectiveness even with limited labeled data.
Table 3: Codebook usage and CER for different sizes of labeled data. The models employ SPIRAL as speech encoder and FSQ as quantizer.
@@ 6 Conclusion
This study proposes the ToneUnit framework, which addresses the “tone shift” issue in synthesized Mandarin Chinese speech based on discrete speech units. By leveraging text data annotated with tone labels as CTC supervision, ToneUnit is able to produce discrete speech units with better phonemic tone discrimination in Mandarin Chinese. Experimental results demonstrate that speech units generated by ToneUnit can synthesize intelligible and natural speech for both Mandarin Chinese and English. Additionally, FSQ is a more effective quantization method to produce tone-aware speech units than VQ, with a much simpler structure. Moreover, ToneUnit works with limited amount of annotated data. We presume that ToneUnit can be used as a general speech discretization method to learn context-aware discrete speech representations.

### v2

@@ 5.3 Ablation study
Table 4: Codebook usage and CER for different sizes of labeled data. The models employ SPIRAL as the speech encoder and FSQ as the quantizer.
@@
A con of ToneUnit is the requirement of annotated data with tone labels. To evaluate the required amount of labeled data, we perform ablation under the same configuration as the FSQ model (ID 4) with different sizes of the AISHELL-1 training set. We randomly sample 3 training subsets of 1, 10, and 50 hours from the full 150-hour training set for the fine-tuning process. We still train the speech synthesizer with the full 150-hour AISHELL-1 dataset for 100k updates. Note that the speech synthesizer is trained with extracted discrete speech units without any text labels. The results are reported in Table 4 in terms of codebook usage and CER. FSQ can still fully use the codewords at different sizes of labeled data. With just 1 hour of annotated data with tone labels, the ToneUnit framework successfully generates speech units capable of synthesizing Mandarin Chinese speech at acceptable machine intelligibility. With a subset of 10-hour training data, the CER of synthesized speech is already comparable to those from the full training set. This suggests that the proposed ToneUnit framework retains effectiveness even with limited labeled data.
@@ 6 Conclusion
This study proposes the ToneUnit framework, which addresses the “tone shift” issue in synthesized Mandarin Chinese speech based on discrete speech units. By leveraging text data annotated with tone labels as CTC supervision, ToneUnit is able to produce discrete speech units with better phonemic tone discrimination in Mandarin Chinese. Experimental results demonstrate that speech units generated by ToneUnit can synthesize intelligible and natural speech for both Mandarin Chinese and English. Additionally, FSQ is a more effective quantization method to produce tone-aware speech units than VQ, with a much simpler structure. Moreover, ToneUnit works with a limited amount of annotated data. We hypothesize that ToneUnit can serve as a connector between the speech modality and the large language model (LLM) backbone in multimodal LLMs. This potential application will be investigated in our future work.